OpenStudy (anonymous):

For each function find f(-2), f(-0.5), and f(3). 7.f(x)= -x +4

OpenStudy (unklerhaukus):

so the function f is a function of the variable x , we write f(x), and say 'f of x' when x is at the point -2 f of x is f(-2) the value of the function f(x) = -x+4 at the point x=-2 is thus f(-2) = -(-2)+4 = [ simplify this ]

OpenStudy (anonymous):

How would I simplify that?

OpenStudy (anonymous):

Would it be f(-2) =6?

OpenStudy (unklerhaukus):

do you understand what i have said? to simplify f(-2) = -(-2) +4 you have to use the fact that the negative of a negative becomes a positive so -(-2) is equal to +2 f(-2) = -(-2) +4 = 2 +4 =

OpenStudy (unklerhaukus):

oh yeah i see you got 6 , thats right!

OpenStudy (unklerhaukus):

now f(-0.5) is the same technique sub x = (-0.5) into f(x) = -(x) +4 f(-0.5) = -(-0.5) +4 and simplify... =

OpenStudy (anonymous):

f(-0.5) =4.5? Could you help me with this also For each function to determine whether y varies directly with x. If so find the constant of variation and write the function. x = -2, 1, 2 y = 3, 4 , 7 and I got 3/-2 = 4,1 = 7/2, but I don't understand, because y doesn't vary directly with x, but it doesn't seem finished.

OpenStudy (unklerhaukus):

f(-0.5) =4.5, yes!

OpenStudy (unklerhaukus):

for this next problem you have these points (x, y) = (-2,3), (1,4), (2,7) right?

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

if you plot the points and try to draw a straight line through them , what do you get ?

OpenStudy (unklerhaukus):

i got a line that missed all the points, or a bendy line,

OpenStudy (anonymous):

Yes, I got the same result. . .

OpenStudy (unklerhaukus):

it can't be a direct variation then

OpenStudy (unklerhaukus):

a direct variation would be a straight line, all the ratios you calculated would have had to be the same

OpenStudy (anonymous):

Oh, so y doesn't vary with x, there is no function rule?

OpenStudy (unklerhaukus):

y does vary with x, but it is not a direct variation, it might be a quadratic , or a higher order curve its hard to tell with only three points

OpenStudy (unklerhaukus):

but you can tell for certain, that if there is a function that does contain those points it will not be of the form y(x) = mx+b

OpenStudy (unklerhaukus):

for y to not vary with x, would be something like (x, y) = (0,4), (1,4), (500,4) here y is constant y(x) = 4

OpenStudy (anonymous):

Okay, what about (1,1), (2,2), (5,5)? I got a strong positive correlation. This would be a direct variation?

OpenStudy (unklerhaukus):

yeah those all fit on a nice straight line, can you tell me what the linear function is that contain those points? y(x) =

OpenStudy (anonymous):

Mmm, don't know. . .y(x)=8?

OpenStudy (unklerhaukus):

(1,1), (2,2), (5,5) y(1) = 1 y(2) = 2 y(5) = 5 y(x) =

OpenStudy (anonymous):

1?

OpenStudy (unklerhaukus):

its a straight line, so it will be of the form y(x) = mx+b

OpenStudy (unklerhaukus):

the line that contains the points (1,1), (2,2), (5,5) is y(x) = x

OpenStudy (anonymous):

y(x) = 5?